OpenStudy (joannablackwelder):

Calculus question: We are given that f(3) = 4, f′(3) = 8, f′′(3) = 11. Use this information to approximate f(3.15).

OpenStudy (joannablackwelder):

Is there a better method of doing this than just using Newton's method and ignoring the second derivative info?

OpenStudy (joannablackwelder):

@IrishBoy123

OpenStudy (irishboy123):

Joanna, i would go with the linear approximations you suggest.

OpenStudy (joannablackwelder):

Thanks! @UnkleRhaukus Any thoughts?

OpenStudy (inkyvoyd):

do a quadratic approximation

OpenStudy (inkyvoyd):

f''(x) will be constant near f(x)

OpenStudy (joannablackwelder):

Oh, I see, I found a formula for that. Thanks so much!

OpenStudy (inkyvoyd):

... and I deleted it because it's not right. Sigh, been so long since I've used calc 2. Anyways, fun fact, the taylor series is basically a continuation of this... https://en.wikipedia.org/wiki/Taylor_series#Calculation_of_Taylor_series

OpenStudy (unklerhaukus):

Assuming \[f(x) = ax^2+bx+c\]\[f'(x) = 2ax+b\]\[f''(x) = 2a\]

OpenStudy (unklerhaukus):

\[f''(3) = 2a =11\implies a=11/2\]

OpenStudy (joannablackwelder):

Using the quadratic approximation that I linked to before, I get about 5.32

OpenStudy (unklerhaukus):

\[f'(3)=8=11(3)+b\implies b=-25\] \[f(3)=4=\tfrac{11}2(3)^2-25(3)+c\implies c=29.5\]

OpenStudy (unklerhaukus):

\[f(x) = \tfrac{11}2x^2-25x+29.5\] \[f(3.15) = \tfrac{11}2(3.15)^2-25(3.15)+29.5 = 5.32375\]

OpenStudy (freckles):

I wonder if taylor series is what they want \[f(x) \approx f(a)+f'(a)(x-a)+\frac{f'(a)}{2}(x-a)^2 \text{ where } a \text{ here equals } 3\]

OpenStudy (freckles):

oh you mentioned that formula in that link @joannablackwelder nevermind

OpenStudy (joannablackwelder):

Thanks so much everybody!